how do solve this using the elimination method 3a+5b=9
4a+3b=1

Answer by mananth(16946)

3.00 a + 5.00 b = 9.00 .............1
Total value
4.00 a + 3.00 b = 1.00 .............2
Eliminate y
multiply (1)by -3.00
Multiply (2) by 5.00
-9.00 a -15.00 b = -27.00
20.00 a + 15.00 b = 5.00
Add the two equations
11.00 a = -22.00
/ 11.00
a = -2.00
plug value of a in (1)
3.00 a + 5.00 b = 9.00
-6.00 + 5.00 b = 9.00
5.00 b = 9.00 + 6.00
5.00 b = 15.00
b = 3.00
a=-2,b=3